OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes though the point (0, 2).

OpenStudy (anonymous):

y = 5x + 2 y = -2/5x + 2 y = 2x + 5 y = 1/2x + 5